# Project Page: https://yongtaoge.github.io/projects/humanwild

### v2

@@ 3.3 Initial Human Image and Annotation Generation
Camera simulation. One drawback of vision-based motion capture systems is that they need to calibrate and synchronize the camera’s intrinsic and extrinsic parameters during the capturing. Thus, the collected human data are limited in terms of the scales and view diversity. On the contrary, our pipeline gets rid of the physical RGBD cameras and can simulate arbitrary human scales and body orientation. Specifically, we randomly determine the orthographic scale $s$ of the human body ($s\in[0.45,1.1]$), along with the horizontal shift ($t_{x},t_{y}$) within a range of $[-0.4/s,0.4/s]$. This methodology ensures that the majority of body parts are visible in the image. Following [27, 77], we determine the translation of the body as $transl=[t_{x},t_{y},f/s]$. The focal length in normalized device coordinate (NDC) space, denoted as $f$, can be computed using the formula $f=1/\tan(FoV/2)$. Here, $FoV$ represents the Horizontal Field of View angle, which is randomly zoomed in from 25 to 120 degrees by following [5, 77].
Image condition generation. To synthesize realistic human images with paired pose annotations, we leverage ControlNet, equipped with the state-of-the-art diffusion model, SDXL [58], as our image generator. Existing ControlNet variants take a 2D skeleton, depth map, or canny map as condition inputs. These inputs are typically detected from real-world images by pre-trained perception models. However, there exist two main drawbacks to generating image conditions from these pre-trained perception models. On one hand, it’s laborious to crawl diverse human pose and shape images from the Internet. On the other, the perception models cannot ensure the generation of fully accurate annotations, thus the different modalities annotations have discrepancies, e.g., the 2D keypoint heatmap from a 2D pose estimator and the depth map from a depth estimator are not aligned. In such cases, if we take the perception results as the multi-condition inputs of ControlNet, the generated images would probably have weird artifacts.
To resolve the problem, we construct the input of ControlNet by taking advantage of the 3D human parametric model, SMPL-X. There exist several large-scale human motion capture databases [44, 5] with diverse body poses and shapes in SMPL-X format. Thanks to the disentanglement of the pose and shape parameters of the SMPL-X model, we can even recombine the two parameters to generate a human mesh that does not exist in the databases. For example, an overweight man doing an extremely difficult yoga pose. Upon getting the simulated camera parameters aforementioned in Sec. 3.3 and 3D human mesh from SMPL-X, we can render an existing 3D human mesh into the image plane, as such, getting the corresponding surce normal map. Notably, the surface normal map is proven to be crucial to generating accurate body shapes and orientation.
@@
Human &\&& scene positioning. We offer an optional flexible background selection feature, allowing users to generate scenes based on text prompts solely or in conjunction with normal maps. To integrate specific backgrounds, we randomly select scene meshes from the ScanNet++ dataset [85] and utilize Octree [48] to partition the room mesh into discrete voxels. With vertex-level classification annotations, we efficiently identify ground plane normals and height to anchor the human mesh. We randomly allocate human mesh within unoccupied areas of the voxel space. Subsequently, we refine human mesh positioning by optimizing the translation using the Chamfer distance to preclude inter-mesh collisions. Finally, we simulate a random camera perspective. The intrinsic matrix is obtained with previous rules, while the extrinsic matrix is crafted using a random azimuth angle. We sample the camera’s height range [-1m, +1m] relative to the pelvis height, and look at a random point on the human torso, to achieve a realistic viewpoint.
@@
Text prompt generation. The aforementioned normal maps have two choices, with or without room mesh as backgrounds. Since our background information is retrieved from Scannet++ [85], for in-the-wild data, we need text prompts to provide background information. Moreover, the human body’s normal maps are not fine-grained enough to determine the gender and appearance of the human. Thus, we incorporate a structured text prompt template to handle this issue. In particular, we designed a simple text template as “A {gender} {action} {environment}". The gender and the action of the person are determined by the SMPL-X annotations. The environment is generated by a large language model, i.e., ChatGPT [51] and LLAMA [70]. To create photo-realistic humans, we also feed negative text prompts, e.g., “ugly, extra limbs, poorly drawn face, poorly drawn hands, poorly drawn feet”, to the model. Finally, we have all of the input conditions of the ControlNet. We apply a total of 40 inference steps for each sample. The resolution of the generated images are all $768\times 768$. The generated images and the input conditions (normal maps, SMPL-X parameters) are regarded as the initial data pairs. See Sup. Mat. for detail analysis.
Figure 4: Data sample visualization of $\tt HumanWild$. For each data sample, the left side is the normal image rendered from SMPL-X model, the right side is the image generated by $\tt HumanWild$ pipeline. The first two rows demonstrate the indoor activities. The third row shows the generated images with huge distortion. The fourth row describes the diverse camera views. The fifth row shows overweight human bodies.
@@ 3.4 Label Denoising
The generated images are not always well-aligned with the input conditions. The most common incorrect case is the generated human and the input conditions form a mirror pair. To resolve this problem, we employ an off-the-shelf foundation segmentation model, i.e., SAM [30], to filter negative samples from the final dataset. Specifically, we first calculate the ground-truth mesh segmentation mask by rendering the 3D mesh into the image plane. Then, we sample a random point coordinate from the ground-truth segmentation mask and feed the generated image and this point coordinate into the SAM model, thereby we can get the predicted human mask for the generated image. Finally, we compute the intersection-over-union (IoU) metric between the ground-truth mask and prediction mask. We filter data samples with IoU lower than $0.8$.
@@ 4.1 Implement Details and Evaluation Metrics
To illustrate the effectiveness and efficiency of our proposed $\tt HumanWild$, we report Frechet Inception Distance (FID [18]) and Kernel Inception Distance (KID [4]), which are widely used to measure the quality of the synthesized images. We measure the alignment accuracy between the SMPLX-rendered mask and the generated images with mean intersection over union (mIoU). We use standard metrics to evaluate body pose and shape accuracy. PVE and MPJPE represent the average error in vertices and joints positions, respectively, after aligning the pelvis. PA-MPJPE further aligns the rotation and scale before computing distance. PVE-T-SC is a per-vertex error in a neutral pose (T-pose) after scale correction. For a fair comparison with the CG-rendered counterpart, we sample SMPL-X parameters from the BEDLAM [5] dataset and render them into surface normal maps with random camera parameters illustrated in Sec. 3. We re-implemented a regression-based HPS method, CLIFF [39], for evaluating the effectiveness of the synthesized dataset on a variety of evaluation benchmarks, 3DPW [46], AGORA [55], EgoBody [89], RICH [20], and SSP-3D [65] (for shape evaluation).
@@ 4.2 Ablation on ControlNet Input Conditions
Figure 5: Surface normal visualization. (a), (d) are two images sampled from LAION5B [64]. (b), (e) are estimated by our method. (c), (f) are estimated by OmniData.
In the initial stages of our research, we employed various input conditions, including 2D keypoint heatmaps, depth images, and normal images, to generate data samples. However, a notable challenge emerged with regard to the ambiguity inherent in both the keypoint and depth conditions. Specifically, the diffusion model struggled to accurately discern the front and back of individuals depicted in these data modalities and had difficulty in aligning generated hands and faces with ground-truths (See Sup. Mat. for some failure cases), yet we found surface normal conditional can greatly reduce this kind of ambiguity. Thus, we choose it as the input condition of the ControlNet. To get accurate surface normal maps in the wild, we train a customized diffusion model based on a photo-realistic synthetic indoor dataset, Hypersim [59]. As shown in Tab. 3, our surface normal estimator outperforms OmniData [12] trained on a mixture of datasets, including Hypersim [59] dataset. We also show qualitative surface normal results in  Fig. 5. The results show that our diffusion-based estimator can generate more detailed normal maps when compared with OmniData [12] in real-world scenes. We annotate around 1M image/surface normal pairs from LAION5B [64] and COYO-700M [7] for training ControlNet. Please refer to the supplementary material for more details. Tab. 3 compares our trained ControlNet with two publicly available keypoint-based and depth-based ControlNets [21, 68] in terms of generated image quality and image/ground-truth alignment accuracy. In order to ensure equitable comparison, all ControlNets employ aligned sets of 1,000 input conditions derived from identical SMPL-X parameters and camera specifications. The results demonstrate that our ControlNet consistently outperforms the keypoint-based and depth-based ControlNet.
Table 2: Ablation of different input conditions of ControlNet.
Table 3: Surface normal evaluation on Hypersim dataset.
@@
[…]
@@
OmniData [83] DPT mix 25.3 16.2 30.4
Ours UNet Hypersim 18.9 9.5 23.2
Table 3: Surface normal evaluation on Hypersim dataset.
4.3 Verification of the Necessity of In-the-wild 3D Data
Proper initialization of backbone weights is crucial for the performance of HPS models. Different weight initialization strategies profoundly impact a HPS model’s convergence speed and final accuracy. HMR-benchmarks [54] conduct systematic analysis by initializing the same backbone pre-trained with different sources, i.e., ImageNet classification weights, 2D pose estimation weights from MPII and COCO. The first three columns of Tab. 4 show the result of their paper. In conclusion, employing a backbone pre-trained on in-the-wild pose estimation datasets proves to be beneficial for HPS tasks. Furthermore, the selection of the pose estimation dataset for pretraining significantly influences model performance. Notably, datasets like COCO demonstrate superior performance compared to MPII, likely due to the greater diversity and complexity of wild scenes contained within COCO. Columns 4 to 6 of the Tab. 4 elucidate the impact of backbone pretraining on CG-rendered datasets, BEDLAM [5] and AGORA [55]. A consistent conclusion is drawn from these results that pretraining with 2D wild data also has a great positive impact on CG-rendered datasets. We posit that neither indoor-mocap datasets nor CG-rendered datasets offer a sufficiently varied array of photo-realistic, in-the-wild scenes. Consequently, the network may struggle to acquire robust features conducive to generalizable testing. In the final column of  Tab. 4, the inclusion of in-the-wild synthesized 3D data generated by $\tt HumanWild$ significantly enhances the HPS performance of the CG-rendered datasets. Hence, data synthesized by diffusion models, featuring diverse human identities and real-world scenes, complements CG-rendered data effectively.
Table 4: Ablation experiments on 3D pose and shape estimation. ‘R’ denotes mixed real-world datasets, ‘H’ denotes $\tt HumanWild$, ‘B’ denotes BEDLAM and ‘A’ denotes AGORA. PA-MPJPE, MPJPE and PVE are evaluated on 3DPW. PVE-T-SC is evaluated on SSP-3D.
@@ 4.3 Verification of the Necessity of In-the-wild 3D Data
4.4 Scale-up
Table 5: Scale-up ablation. We study the scaling law of the amount of data and the model sizes. The metrics are MPJPE for Human36M [22] and PVE for other evaluation benchmarks. Foundation models are named “ViT-M", where M indicates the size of ViT backbone (S, B, L, H). MPE: mean primary error cross 5 evaluation benchmarks. AGORA uses the validation set, and EgoBody uses the EgoSet.
@@ 4.4 Scale-up
100% 0.79M ViT-S 23M 106.8 97.3 107.7 106.5 87.1 101.1
100% 0.79M ViT-B 88M 103.6 94.2 105.1 103.5 82.5 97.8
100% 0.79M ViT-L 305M 97.5 85.6 97.2 97.4 75.7 90.7
100% 0.79M ViT-H 633M 90.2 82.3 92.1 90.3 72.4 85.5
We perform experiments to show the effectiveness of the generated datasets by studying the scaling law of the amount of data and the model sizes in Tab. 5. Notably, our training set is independent of the evaluation benchmarks in Tab. 5. We show results with various ViT backbones and percentages of the proposed dataset. It is observed that 1). more training data leads to better performance. 2). A larger backbone achieves higher performance on the proposed dataset.

### v3

@@ 3.3 Initial Human Image and Annotation Generation
Camera simulation. One drawback of vision-based motion capture systems is that they need to calibrate and synchronize the camera’s intrinsic and extrinsic parameters during the capture. Thus, the collected human data are limited in terms of the scales and view diversity. On the contrary, our pipeline gets rid of the physical RGBD cameras and can simulate arbitrary human scales and body orientations. Specifically, we randomly determine the orthographic scale $s$ of the human body, with $s\in[0.45,1.1]$. We also compute the horizontal shifts $t_{x}$ and $t_{y}$ within the ranges $b_{t}/s,b_{b}/s]$ and $b_{l}/s,b_{r}/s]$ respectively. Here, $b_{t},b_{b},b_{l}$, and $b_{r}$ represent the empirical boundary values chosen based on whether a normal map is being generated for the half-body or the whole body. Following [1, 56], we determine the translation of the body as $transl=[t_{x},t_{y},f/s]$. The focal length in normalized device coordinate (NDC) space, denoted as $f$, can be computed using the formula $f=1/\tan(FoV/2)$. Here, $FoV$ represents the Horizontal Field of View angle, which is randomly adjusted from 25 to 120 degrees by following [19, 56]. We also augment the human body’s rotation by $180\degree,180\degree]$ along $y$ axis and $30\degree,30\degree]$ along $x,z$ axis.
Image condition generation. To synthesize realistic human images with paired pose annotations, we leverage the multi-condition ControlNet fine-tuned in Sec.˜3.2, as our image generator. Existing ControlNet variants usually take a 2D keypoint heatmap, depth map, or surface normal map as condition inputs. These inputs are typically detected from real-world images by pre-trained perception models. Different from the common usage, we construct the input of ControlNet by taking advantage of the 3D human parametric model, SMPL-X. Specifically, we use existing large-scale human motion capture databases [27] with diverse body poses and shapes in SMPL-X format to render the spatial input condition maps, i.e., keypoint heatmaps, depth maps, and surface normal maps. Thanks to the disentanglement of pose and shape parameters in the SMPL-X model, we can even recombine these parameters to generate a human mesh that doesn’t exist in the databases.
Upon getting the simulated camera parameters aforementioned in Sec.˜3.3 and 3D human mesh from SMPL-X, we can render an existing 3D human mesh into the image plane by using perspective projection, as such, getting the corresponding source normal map, depth map, and 2D keypoint heatmap. Notably, the surface normal map is proven to be crucial to generating accurate body shapes and orientation, and the keypoint heatmap and depth map are useful for improving the alignment between the input body pose and the generated image.
@@
Human &\& scene positioning. We offer an optional flexible background selection feature, allowing users to generate scenes based on text prompts solely or in conjunction with background-free keypoint heatmaps, depth maps, and normal maps. To integrate specific backgrounds, we randomly select scene meshes from the ScanNet++ dataset [57]. We first utilize Octree [58] to partition the room mesh into discrete voxels. With vertex-level classification annotations, we efficiently identify ground plane normals and height to anchor the human mesh. We randomly allocate human mesh within unoccupied areas of the voxel space. Subsequently, we refine human mesh positioning by optimizing the translation using the Chamfer distance to preclude inter-mesh collisions. Finally, we simulate a random camera perspective. The intrinsic matrix is obtained with the previous rules, while the extrinsic matrix is crafted using a random azimuth angle. We sample the camera’s height range [-1m, +1m] relative to the pelvis height, and look at a random point on the human torso, to achieve a realistic viewpoint. The visualization results can be seen in the appendix.
@@
Text prompt generation. To create controllable and diverse in-the-wild background scenes, we can use text prompts to provide background information. Moreover, the human body’s spatial maps are not fine-grained enough to determine the gender and appearance of a human. Thus, we incorporate a structured text prompt template to handle this issue. In particular, we designed a simple text template as “A {gender} {action} {environment}". The gender of the person is determined by the SMPL-X annotations. The environment is generated by a large language model, i.e., ChatGPT [59] and LLAMA [60]. To create photo-realistic humans, we also feed negative text prompts, e.g., “ugly, extra limbs, poorly drawn face, poorly drawn hands, poorly drawn feet”, to the model.
Finally, we have all of the input conditions of the multi-condition ControlNet. We apply a total of 40 inference steps for each sample. The generated images are up to $1024\times 1024$ resolution. The generated images and the input conditions (SMPL-X parameters) are regarded as the initial data pairs.
@@ 3.4 Label Denoising
Figure 6: Analysis of Mirror Image Generation from Depth Maps: This visualization demonstrates the correction of a common failure mode. The leftmost image is the guiding depth map. The middle image is an incorrect output where the model mistakenly generates the person’s back. The rightmost image shows the desired outcome, a correctly rendered frontal view.
The generated images are not always well-aligned with the input conditions. For example, an incorrect case is that the generated human and the input conditions form a mirror pair, as shown in Fig.˜6. To resolve this problem, we employ an off-the-shelf foundation segmentation model, i.e., SAM [29], to filter negative samples from the final dataset. Specifically, we first calculate the ground-truth mesh segmentation mask by rendering the 3D mesh into the image plane. Then, we sample a random point coordinate from the ground-truth segmentation mask and feed the generated image and this point coordinate into the SAM model, thereby we can get the predicted human mask for the generated image. Finally, we compute the intersection-over-union (IoU) metric between the ground-truth mask and the prediction mask. We filter data samples with IoU lower than $0.8$.
To further improve the alignment between the generated human images and SMPL-X pose parameters, we train a sparse 2D surface estimator by using RTMPose [61] architecture to estimate a total of 128 sparse 2D vertices sampled from dense SMPL-X vertices [62]. The input resolution of the detector is $512\times 384$. As our generated images have accurate camera intrinsics, we can project the ground-truth 3D SMPL-X vertices into 2D vertices using perspective projection. Then, we can optimize the original SMPL-X pose parameters with SMPLify [16]. Note that we only optimize the body pose parameters for 20 iterations and make the shape parameters fixed.
@@ 4 Experiments
4.1 Human-centered Datasets for training ControlNet
In this section, we list our datasets for fine-tuning the multi-condition ControlNet in terms of the number of subjects, the number of scenes, and the details of diversity. Overall, we use three datasets for fine-tuning an existing variant. The RenderedHuman dataset contains 202,667 human images, distinguishing itself by providing an unparalleled wealth of precise annotations, including texts, keypoints, depth, and surface maps. The PexHuman is collected with more than 120 daily human activities, including 100,880 images. We also filter the Laion5B dataset by a human detector, YOLOS [51], where only images with fewer than 3 humans are kept. We generate keypoints, depth, and surface maps for PexelHuman and LaionHuman with DWPose [54], DepthAnythingV2 [63], and DSINE [53], separately. To get the text description of the human images, we leverage a multi-modality vision-language model, CogVLM [64], to generate detailed image captions.
Dataset Data Type # Samples # Scenes # Identies
LaionHuman real 233,728 > 200K > 200K
PexelHuman real 100,880 > 100K > 100K
RenderHuman Synthetic 202,667 660 630
TABLE II: An overview of all datasets for fine-tuning multi-condition ControlNet.
@@ 4.2 Implement Details and Evaluation Metrics
To illustrate the effectiveness and efficiency of our proposed HumanWild, we report the Frechet Inception Distance (FID [65]) and the Kernel Inception Distance (KID [66]), which are widely used to measure the quality of the synthesized images. We measure the alignment accuracy between the SMPLX-rendered mask and the generated images with mean intersection over union (mIoU). We use standard metrics to evaluate body pose and shape accuracy. PVE and MPJPE represent the average error in vertices and joints positions, respectively, after aligning the pelvis. PA-MPJPE further aligns the rotation and scale before computing distance. PVE-T-SC is a per-vertex error in a neutral pose (T-pose) after scale correction. For a fair comparison with the counterpart pure CG pipeline, we sample SMPL-X pose and shape parameters from the AMASS [27] dataset and render them into surface normal maps with random camera parameters illustrated in Sec.˜3. We re-implemented a regression-based HPS method, CLIFF [4], for evaluating the effectiveness of the synthesized dataset on a variety of evaluation benchmarks, 3DPW [23], AGORA [32], EgoBody [67], RICH [68], and SSP-3D [69] (for shape evaluation).
Unless specified, all HPS models are trained on the CLIFF [4] model with an HRNet-48 [70] backbone. Following BEDLAM [19], the supervision losses are a combination of MSE loss on model parameters, projected keypoints, 3D joints, and an L1 loss on 3D vertices.
@@ 4.3 Ablation on ControlNet Input Conditions
In the initial stages of our research, we employed various input conditions, including 2D keypoint heatmaps, depth images, and normal images, to generate data samples. However, a notable challenge emerged with regard to the ambiguity inherent in both the keypoint and depth conditions. Specifically, the diffusion model struggled to accurately discern the front and back of individuals depicted in these data modalities and had difficulty in aligning generated hands and faces with ground-truths (See Sup. Mat. for some failure cases), yet we found that surface normal conditioning can greatly reduce this kind of ambiguity. Thus, we choos a multi-condition variant of the ControlNet, which can optionally treat keypoint, depth, and surface normal maps as input. Table˜III compares our trained ControlNet with two publicly available keypoint-based and depth-based ControlNets [63, 71] in terms of generated image quality and image/ground-truth alignment accuracy. In order to ensure equitable comparison, all ControlNets employ aligned sets of 1,000 input conditions derived from identical SMPL-X parameters and camera specifications. The results demonstrate that our ControlNet consistently outperforms the keypoint-based and depth-based ControlNet.
@@
Condition FID $\downarrow$ $\text{KID}\downarrow$ mIoU $\uparrow$
@@
All conditions 22.5 2.74 59.2
TABLE III: Ablation of different input conditions of ControlNet.
We also visualize the generated images by a general depth-conditioned ControlNet, and our finetuned multi-condition ControlNet in Fig.˜7. As we can see, the general depth-conditioned ControlNet trained on general datasets fails to generate accurate fingers, poses, and limited background scenes, while our fine-tuned multi-condition ControlNet can generate diverse photo-realistic humans tightly aligned with the input condition.
Figure 7: Visualization of a general depth-conditioned ControlNet, and our finetuned multi-condition ControlNet. The first column is the input depth condition. The second column is the prediction from the general depth-conditioned ControlNet. The third column is our v1 model. The fourth column is our current v2 model.
4.4 Verification of the Necessity of In-the-wild 3D Data
Previous studies, HMR-Benchmark [72] and BEDLAM [19] point out that different backbone initialization strategies profoundly impact an HPS model’s convergence speed and final accuracy. HMR-benchmarks [72] conduct a systematic analysis by initializing the same backbone pre-trained with different sources, i.e., ImageNet classification weights, 2D pose estimation weights from MPII, and COCO. As shown in the first three columns of Table˜VI show the results are trained on real datasets. The conclusion is that COCO keypoint pretraining helps HPS models trained on real datasets achieve the best performance.
To verify whether the conclusion still holds for CG-rendered synthetic datasets, columns 4 to 6 of the Table˜VI elucidate the impact of backbone pretraining on BEDLAM [19] and AGORA [32]. A consistent conclusion is drawn that COCO keypoint pretraining also has a great positive impact on CG-rendered datasets.
Based on the above observation, we posit that neither indoor-mocap datasets nor CG-rendered datasets offer a sufficiently varied array of photo-realistic, in-the-wild scenes. Consequently, the HPS network, which is not initialized with large-scale in-the-wild data pretraining, struggles to perform well on unseen scenes and has limited generalization capability. As our generated data contains diverse in-the-wild scenes, we add the data generated by our pipeline for training HPS models. As we can see in the final column of  Table˜VI, the inclusion of in-the-wild synthesized data generated by HumanWild can further enhance the HPS performance of the CG-rendered datasets. Hence, data synthesized by diffusion models can effectively complement CG-rendered data by featuring diverse human identities and real-world scenes.
@@ 4.4 Verification of the Necessity of In-the-wild 3D Data
PDHuman 102.4 159.1 168.0
HumanWild 90.0 151.8 160.9
TABLE IV: HPS results on SPEC-MTP.
Training Data MPJPE $\downarrow$ PA-MPJPE $\downarrow$ PVE $\downarrow$ PA-PVE $\downarrow$
B+A+P 143.1 94.9 163.9 95.7
B+A+P+H 134.6 89.1 154.1 88.9
TABLE V: HPS results on MOYO.
@@
TABLE VI: Ablation experiments on 3D pose and shape estimation. ‘R’ denotes mixed real-world datasets, ‘H’ denotes HumanWild, ‘B’ denotes BEDLAM and ‘A’ denotes AGORA. PA-MPJPE, MPJPE and PVE are evaluated on 3DPW. PVE-T-SC is evaluated on SSP-3D.
4.5 Data Scale-up Ablation Study
@@
100% 0.63M ViT-S 23M 106.8 97.3 107.7 106.5 87.1 101.1
100% 0.63M ViT-B 88M 103.6 94.2 105.1 103.5 82.5 97.8
100% 0.63M ViT-L 305M 97.5 85.6 97.2 97.4 75.7 90.7
100% 0.63M ViT-H 633M 90.2 82.3 92.1 90.3 72.4 85.5
TABLE VII: Scale-up ablation. We study the scaling law of the amount of data and the model sizes. The metrics are MPJPE for Human36M [12] and PVE for other evaluation benchmarks. Foundation models are named “ViT-M", where M indicates the size of ViT backbone (S, B, L, H). MPE: mean primary error cross 5 evaluation benchmarks. AGORA uses the validation set, and EgoBody uses the EgoSet.
We perform experiments to show the effectiveness of the generated datasets by studying the scaling law of the amount of data and the model sizes in Table˜VII. Notably, our training set is independent of the evaluation benchmarks in Table˜VII. We show results with various ViT backbones and percentages of the proposed dataset. It is observed that 1). Scale-up synthetic training data generated by HumanWild leads to better performance. 2). A larger backbone benifits more from a larger synthetic training dataset.